OpenStudy (anonymous):

Find the sin,cos, tan of the follwing angle pi/6

jimthompson5910 (jim_thompson5910):

use this unit circle http://etc.usf.edu/clipart/43200/43215/unit-circle7_43215_lg.gif locate the angle pi/6 then mark the corresponding point to this angle the x coordinate of this point is equal to `cos(pi/6)` the y coordinate of this point is equal to `sin(pi/6)` divide `sin(pi/6)` over `cos(pi/6)` to get `tan(pi/6)`

OpenStudy (anonymous):

do i get 3 answer?

jimthompson5910 (jim_thompson5910):

yes you'll have 3 different results for sine, cosine, tangent

OpenStudy (anonymous):

DO i write the answer in degree or fraction?

jimthompson5910 (jim_thompson5910):

they'll be fractions with square roots in them

jimthompson5910 (jim_thompson5910):

well except for sin(pi/6)

OpenStudy (anonymous):

can u show me the exampple of the answer im confused what should i write

jimthompson5910 (jim_thompson5910):

see attached

jimthompson5910 (jim_thompson5910):

let me know if that helps or not

OpenStudy (anonymous):

but do i have to write 3 answer?

jimthompson5910 (jim_thompson5910):

do you see how I got \[\Large \cos\left(\frac{\pi}{6}\right) = \frac{\sqrt{3}}{2}\] ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

im so confuse

jimthompson5910 (jim_thompson5910):

whats the y coordinate of that point

OpenStudy (anonymous):

is the answer just radical 3/2

jimthompson5910 (jim_thompson5910):

that's for cos(pi/6)

OpenStudy (anonymous):

what bout sin or tan